retierment funds
five and a half years ago.chris invested 10,000in a reterminent fund that grew at thr rate of 10.82%/year compounded Quarterly what is the account worth today

Answer by rfer(16322)

A=10000(1+.1082/4)^4*5.5
A=10000(1.02705)^22
A=10000*1.798933
A=17,989.33